OpenStudy (anonymous):

Need help fraction arithmetic 1. 6/7 x 5/3 2. 4 2/5 x (-3 1/3) 3. 2 2/5 divide (-1 1/2) 4. - 8/9 divide (- 4/3) 5. 5/8 x (- 4/15) word problem 6. My recipe calls for 2/3 cups of white flour and 2 1/2 cups of whole wheat flour. How much flour do i need in total for my recipe?

OpenStudy (anonymous):

Can anyone help me step by step for all of them

OpenStudy (anonymous):

hey can u help

OpenStudy (anonymous):

1.\[6/7 * 5/3 = (3*2)/7 * 5/3 = 2/7 * 5 = 10/7\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

where do u get the 2 for number 1 ?

OpenStudy (anonymous):

2. \[42/5 * (-31/3) = (3*14) * (-31/3) = 14/5 * (-31) = -434/5\]

OpenStudy (anonymous):

kk im just going to listen to u

OpenStudy (anonymous):

6 = 3 * 2 right?

OpenStudy (anonymous):

we simplify 3 in the first nominator with 3 from the second denominator

OpenStudy (anonymous):

clear?

OpenStudy (anonymous):

clear